The empirical formula of styrene is CH; its molar mass is 104.1 g/mol. What is the molecular formula
1
Expert's answer
2020-05-03T15:09:47-0400

(Empirical formula)×n=Molecular formula

(CH)×n=104.1 g/mol

(12.01 g/mol + 1.01 g/mol)×n=104.1 g/mol

(13.02 g/mol)×n=104.1 g/mol

n=7.99

Answer: The molecular formula is C8H8